OpenStudy (anonymous):

Will fan and medal please help! you owe $1568 on a credit card with a limit of 2,200 at an interest rate of 11.3 APR. you pay 300 a month until it is paid off. how may months does it take to pay it off and how much interest did you pay? I am having trouble with the interest any help?

OpenStudy (mathmate):

Assuming you make the first payment at the end of the month. Let R = 1+0.113/12 = 1.009416667 = 1 + monthly interest Interest is assumed to be compounded every month. The amount owing at the end of the first month = 1568*1.00941667(owing+interest)-300 (payment) = 1282.77 The amount owing at the end of the second month = 1282.77*1.00941667(owing+interest)-300 (payment) = 994.84 Continuing this way, you will get the amount owing decreasing to 704.21 410.844 114.71 -184.21 (after the sixth payment). Therefore he has paid 1800-184.21=$1615.79 instead of $1568 that he owed initially. Do you follow how the solution works? @saysabr

OpenStudy (anonymous):

so, that is another account of mine, and im not following completely...so it takes 6 months by paying 300 a month and the interest payed was 47.79?

OpenStudy (anonymous):

@mathmate

OpenStudy (mathmate):

Yes, that's correct. For the last month, he does not pay 300, but 300-184.21=$115.79. The interest is 1615.79-1568=$47.79 over the six months. Approximately it is 47.79/1568=3.04% which is the right ball park figure. (he owes approximately half of 1568 for six-months, so about 1/4 of the annual interest).